OpenStudy (thivitaa):

if x^2+y^2=2y.. show that d^2y/dx^2=1/(1-y)^3

OpenStudy (anonymous):

something seems wrong with this

OpenStudy (anonymous):

As given :- x^2 + y^2 = 2y Therefore y = (x^2 + y^2) / 2 Hence d^2 (y) / dx^2 = dy/dx * dy/dx = (dy/dx)^2 dy/dx = d [(x^2 + y^2) / 2] / dx = d(x^2/2) + d(y^2 /2) / dx = (1/2) (2x) + (1/2) (2y) = x + y Therefore ( dy/dx )^2 = x^2 + y^2 + 2xy Here on replacing the value of x from the given eqaution we get the answer

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/4130/if-x-2-y-2-2y-show-that ask your calculus homework questions at http://www.mathskey.com/question2answer/calculus … and get free math help. all the best